If (A/4sin15degrees^2)+(A/4cos15degrees^2)=8 the value of A is?
Answer by lwsshak3(11628)

If (A/4sin15degrees^2)+(A/4cos15degrees^2)=8 the value of A is?
(A/4)(sin^2(15)+cos^2(15)=8
A/4*(1)=8
A=32